OpenStudy (anonymous):

Take -4 + 3a^(2) from 7a - a^(2) -4a^(2)- 7a- 4 -4a^(2)- 7a+ 4 -4a^(2)+ 7a+ 4 4a^(2)+ 7a+ 4

OpenStudy (anonymous):

maybe i could help you...\[-(4a^2+7a+4)\] dividing by -1... \[4a^2+7a+4\] now... \[\frac{ -7\pm \sqrt{7^2-4\times(4\times4}) }{ 2\times4 }\]

OpenStudy (anonymous):

i bet you can do it now! okay? @mabanana